let: y=14+2x1+x2-0.12x1-0.08x2+0.12x1x2
given: Py= $5/unit
Px1=$1/unit
Px2=$2/unit
Fixed cost= $200
find
a. the amount of (x1) and (x2) that maximizes profit
b. the amount of (y) at maximum profits
c. the amount of profit
d. the maximum amount of (y) that can be produced
Answer #1

Cajeta Given Criven Y= 14+ 24 +22*0.1221-0.882 + 0*122492 Py = $5/unit ; free Pay = $ / unit 29. 3) - P2 - 42 unit now, MaxWiese I The Amount of y at Maximum profit - X ya [14+7881.69), 4.0-26-4), +0.12(1.67)(-4) 214+ 3.13 – 2.8 – 2.80 = 12.13- 2.